OpenStudy (anonymous):

MEDAL FOR THE CORRECT ANSWER! Given f(x) = the quantity of 5x+1/2 solve for f-1(8). 3 7 10 20

OpenStudy (campbell_st):

so is it \[f(x) = \frac{5x +1}{2}\]

OpenStudy (anonymous):

@campbell_st Yes :)

OpenStudy (campbell_st):

ok... to find the inverse let the original equation be \[y = \frac{5x + 1}{2}\] to find the inverse swap x and y so \[x = \frac{5y + 1}{2} \] now make y the subject of the equation.... this will give the inverse

OpenStudy (anonymous):

@campbell_st first option?

OpenStudy (campbell_st):

it is